# A Compact Post-quantum Strong Designated Verifier Signature Scheme from Isogenies

### v1

@@ Abstract
Digital signatures are essential cryptographic tools that provide authentication and integrity in digital communications. However, privacy-sensitive applications—such as e-voting and digital cash—require more restrictive verification models to ensure confidentiality and control. Strong Designated Verifier Signature (SDVS) schemes address this need by enabling the signer to designate a specific verifier, ensuring that only this party can validate the signature. Existing SDVS constructions are primarily based on number-theoretic assumptions and are therefore vulnerable to quantum attacks. Although post-quantum alternatives—particularly those based on lattices—have been proposed, they often entail large key and signature sizes.
In this work, we introduce CSI-SDVS, a novel isogeny-based SDVS scheme that offers a compact, quantum-resistant alternative. Our construction builds on the ideal class group action framework of CSIDH and the signature techniques of CSI-FiSh, and relies on the hardness of the Multi-Target Group Action Inverse Problem (MT-GAIP). CSI-SDVS achieves strong security guarantees—namely, Strong Unforgeability under Chosen-Message Attacks (SUF-CMA), Non-Transferability (NT), and Privacy of Signer’s Identity (PSI)—in the random oracle model. Remarkably, both the keys and signatures in CSI-SDVS are of size $\mathcal{O}(\lambda)$, representing a significant improvement over the typical $\mathcal{O}(\lambda^{2})$ bounds in existing post-quantum SDVS schemes, thereby making it among the most compact PQC-based SDVS schemes and the only post-quantum secure construction based on isogenies.
@@ 1.1 Related Work
Post-Quantum Constructions. In response to the vulnerabilities of classical assumptions, post-quantum SDVS schemes have been developed based on quantum-resistant foundations. Code-based SDVS constructions were proposed by Assar [8], Ren et al. [9], and Shooshtari et al. [10], leveraging the hardness of the bounded syndrome decoding problem in coding theory [11]. While these schemes are believed to be quantum-resistant, it was shown in [12] that the constructions by Ren et al. and Shooshtari et al. fail to satisfy the critical non-transferability property, which is central to the SDVS notion.
One of the earliest lattice-based SDVS constructions was introduced by Wang et al. [14], employing preimage sampleable functions (PSFs) and Bonsai Trees for trapdoor delegation. Their work builds on the influential framework of Gentry et al. [13], which introduced PSFs as a core primitive in lattice-based signatures. Although the construction offers provable security in the random oracle model, it suffers from efficiency issues due to large key and signature sizes. Later, in 2016, Noh et al. [15] proposed a scheme in the standard model based on an LWE-based encryption scheme and a lattice-based chameleon hash function, albeit at the cost of significant computational overhead. To improve practicality, Cai et al. [16] introduced a more efficient scheme based on the Ring-SIS assumption, using rejection sampling to reduce signature size while maintaining security. More recently, Zhang et al. [17] proposed an SDVS construction combining SIS and LWE assumptions, further contributing to the development of quantum-secure SDVS protocols. However, lattice-based schemes generally incur key and signature sizes with complexity $\mathcal{O}(\lambda^{2})$ [14, 15, 16, 17], which may limit their suitability for resource-constrained environments.
@@ 1.2 Our Contribution
  • We present CSI-SDVS, an efficient post-quantum SDVS scheme constructed via class group actions on the set of $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves. The security of our scheme is based on the hardness of the Multi-Target Group Action Inverse Problem (MT-GAIP) [24], and the design draws inspiration from the CSIDH [23] and CSI-FiSh [24] frameworks.
@@
  • In terms of efficiency, our construction achieves key and signature sizes of asymptotic complexity $\mathcal{O}(\lambda)$, significantly improving over the $\mathcal{O}(\lambda^{2})$ bounds common in existing lattice-based SDVS schemes [14, 15, 16, 17].
@@
Outline.
In the next section, we recall the formal definition of an SDVS scheme and provide the associated security requirements that such a scheme must satisfy. We then present a brief overview of the necessary mathematical prerequisites and describe the underlying group action-based hardness assumptions upon which our construction relies. In Section 3, we introduce our proposed SDVS scheme, CSI-SDVS. Section 4 analyzes the efficiency of our construction, and Section 5 provides a formal security proof.
@@ 2.1 Strong Designated Verifier Signature (SDVS) Scheme
An SDVS scheme involves two distinct parties: a signer $\mathsf{S}$ and a designated verifier $\mathsf{V}$, and is defined over a message space $\mathcal{M}$ and a signature space $\mathsf{\Sigma}$. In such a scheme, the signer $\mathsf{S}$, holding a secret key $\mathsf{sk_{S}}$, generates a signature $\sigma\in\mathsf{\Sigma}$ on a message $m\in\mathcal{M}$ using its secret key and the public key of the designated verifier. The resulting signature $\sigma$ can be verified only by $\mathsf{V}$ using its secret key $\mathsf{sk_{V}}$, ensuring that no third party can verify the signature or be convinced of its validity. This restriction preserves the signer’s privacy by eliminating public verifiability while maintaining correctness for the designated verifier. Moreover, the designated verifier is equipped with a simulation algorithm $\mathsf{Simul}$ that can generate a simulated signature $\sigma’$ ’ on any message $m’\in\mathcal{M}$ using $\mathsf{sk_{V}}$ and the signer’s public key, thereby preventing any verifiable evidence of the signer’s involvement. The formal syntax of an SDVS scheme is presented below, following the framework introduced in [18].
@@ 2.2 Security Model
Query Phase: In this phase, $\mathcal{C}$ responds to polynomially many adaptive queries made by $\mathcal{A}$ by following the steps described below: Signing Oracle $\mathcal{O}_{\mathsf{Sign}}$: On receiving queries on a message $m$ from $\mathcal{A}$, $\mathcal{C}$ runs $\textsf{Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m)$ to obtain a signature $\sigma$ and sends it to $\mathcal{A}$. It then sets $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$. Simulating Oracle $\mathcal{O}_{\mathsf{Simul}}$: On receiving queries on a message $m$ from $\mathcal{A}$, $\mathcal{C}$ runs $\textsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m)$ to obtain a signature $\sigma$ and sends it to $\mathcal{A}$. It then sets $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$. Verifying Oracle $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving a message-signature pair $(m,\sigma)$ adaptively from $\mathcal{A}$, $\mathcal{C}$ runs $\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},\sigma,m)$ to obtain a decisional value $1$ for valid and 00 otherwise, and sends it to $\mathcal{A}$. Output: Eventually, $\mathcal{A}$ outputs a forgery $(m^{\ast},\sigma^{\ast})$, and wins if the following conditions hold: $\textsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},\sigma^{\ast},m^{\ast})=1.$ $(m^{\ast},\sigma^{\ast})\not\in\mathcal{Q}.$
@@
In an SDVS scheme, the notion of non-transferability (NT) ensures that the designated verifier cannot convince any third party of the authenticity of a valid signature. This is achieved by enabling the verifier to use the Simul algorithm to generate simulated signatures that are computationally indistinguishable from genuine ones produced by the signer. As a result, any conviction about the validity of a signature cannot be reliably transferred to others. The NT security of our SDVS construction is formalized as follows:
@@ 2.3 Elliptic Curves and Isogenies
Let $k:=\mathbb{F}_{q}$ denote a finite field, where $q:=p^{n}$ for a prime $p>3$ and a positive integer $n$. An elliptic curve $E$ over $k$ is a smooth, projective, genus-1 curve defined over $k$ with a distinguished point $\mathcal{O}_{E}$ serving as the identity element for the group law on $E$. For a positive integer $\ell$, the $\ell$-torsion subgroup of $E$ is defined as $E[\ell]:=\{P\in E(\overline{k})\mid[\ell]P=\mathcal{O}_{E}\}$. An elliptic curve is called supersingular if its group of $p$-torsion points is trivial over $\overline{\mathbb{F}}_{p}$, i.e., $E[p]=\{\mathcal{O}_{E}\}$. In particular, if $E/\mathbb{F}_{p}$ is supersingular, then #$\#E(\mathbb{F}_{p})=p+1$.
@@
An endomorphism is an isogeny from an elliptic curve $E$ to itself. Examples include the multiplication-by-$m$ map $[m]:P\mapsto[m]P$ for $m\in\mathbb{Z}$, and the Frobenius endomorphism $\pi:(x,y)\mapsto(x^{q},y^{q})$ for an elliptic curve defined over $\mathbb{F}_{q}$. The set of all endomorphisms of $E$ forms a ring under addition and composition, known as the endomorphism ring of $E$ and denoted by $\mathrm{End}(E)$. For a supersingular elliptic curve $E$ defined over $\mathbb{F}_{p^{2}}$, the $\mathrm{End}(E)$ is isomorphic to a maximal order in a quaternion algebra that is ramified at $p$ and $\infty$. In contrast, for a supersingular elliptic curve defined over $\mathbb{F}_{p}$, the ring of $\mathbb{F}_{p}$-rational endomorphisms, denoted by $\mathrm{End}_{\mathbb{F}_{p}}(E)$, is isomorphic to an order in some imaginary quadratic field. Hence, we have a strict inclusion $\mathrm{End}_{\mathbb{F}_{p}}(E)\subset\mathrm{End}(E)$.
@@
Notation. Following [24], if $\mathrm{Cl}(\mathfrak{O})$ is cyclic with generator $\mathfrak{g}$, then the map $\mathbb{Z}/N\mathbb{Z}\xhookrightarrow{}\mathrm{Cl}(\mathfrak{O})$ defined by $a\mapsto\mathfrak{g}^{a}$ gives a convenient representation of class group elements. Here, $N=\#\mathrm{Cl}(\mathfrak{O})$, and any element $\mathfrak{a}\in\mathrm{Cl}(\mathfrak{O})$ can be written as $\mathfrak{g}^{a}$ for some $a\in\mathbb{Z}/N\mathbb{Z}$. Using the shorthand $[a]$ for $\mathfrak{g}^{a}$ and $[a]E:=\mathfrak{g}^{a}*E$, it follows that $[a][b]E=[a+b]E$.
@@ 2.4 Hardness Assumptions
(Group Operations) Given strings $g$ and $h$, decide if they represent elements in $G$ and whether they are equal. Given $g,h\in G$, compute $g^{-1}$, $gh$, and decide if $g=h$.
@@ 3 The Scheme: CSI-SDVS
In this section, we present CSI-SDVS, a new isogeny-based SDVS scheme. The construction integrates the CSIDH framework with the exact group action structure introduced in CSI-FiSh. This hybrid design preserves the post-quantum security guarantees of class group actions. By leveraging the explicitly known cyclic structure of the class group for CSIDH-friendly primes (as shown in [24]), deterministic computation and uniform randomness are naturally enabled. As a result, we achieve efficient and verifiable designated signatures without relying on heuristic assumptions.
In our scheme, CSI-SDVS, the key generation procedures for both the signer and the designated verifier closely follow the key generation phase in CSI-FiSh. At the heart of the $\mathsf{Sign}$ procedure—also mirrored in $\mathsf{Simul}$—lies a structured interaction between class group elements and $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves. Starting from a base curve $E_{0}$, the signer selects a secret randomness $b\in\mathbb{Z}_{N}$ and computes the intermediate curve $\bar{E}=[b]E_{0}$ using the known group action. The designated verifier, holding a secret key $v\in\mathbb{Z}_{N}$, then computes the target curve $Y=[v]\bar{E}=[v+b]E_{0}$, which serves as the core component for verifying the designated verifier signature.
@@
A trusted party executes the setup procedure to generate the public parameters $\mathsf{pp}$ as follows. First, it selects a large prime $p=4\ell_{1}\ell_{2}\dots\ell_{n}-1$, where the $\ell_{i}$ are small, distinct, odd primes. Specifically, it sets $n=74$, $\ell_{1}=3$, $\ell_{73}=373$, and $\ell_{74}=587$. Next, it fixes a base curve $E_{0}:y^{2}=x^{3}+x$ defined over $\mathbb{F}_{p}$, such that $E_{0}\in\mathcal{E}ll_{p}(\mathfrak{O})$, and chooses a generator $\mathfrak{g}$ of the class group $\mathcal{G}=\text{Cl}(\mathfrak{O})$ with class number $N\approx 2^{\lambda}\approx p^{1/2}$, where $\mathfrak{O}=\mathbb{Z}[\sqrt{-p}]$. It then samples a cryptographically secure hash function :$\mathcal{H}:\{0,1\}^{\ast}\rightarrow\{0,1\}^{\lambda}$, where $\lambda$ is a security parameter. In addition, a positive integer $\eta$—which depends on the security parameter—is specified. Finally, it outputs the public parameters $\mathsf{pp}:=(p,\mathfrak{g},N,E_{0},\mathcal{H},\eta)$.
@@
    Compute the curves $Y^{\prime}_{i}=[y_{i}]E_{i}$ for each $i\in[\eta]$.

### v3

@@ Abstract
Digital signatures are fundamental cryptographic tools that provide authentication and integrity in digital communications. However, privacy-sensitive applications—such as e-voting and digital cash—require more restrictive verification models to ensure confidentiality and control. Strong Designated Verifier Signature (SDVS) schemes address this need by enabling the signer to designate a specific verifier, ensuring that only this party can validate the signature. Existing SDVS constructions are primarily based on number-theoretic assumptions and are therefore vulnerable to quantum attacks. Although post-quantum alternatives—particularly those based on lattices—have been proposed, they often entail large key and signature sizes.
In this work, we present $\mathsf{CSI\text{-}SDVS}$, a novel isogeny-based SDVS scheme that offers a compact, quantum-resistant alternative to existing SDVS constructions. The scheme leverages the ideal class group action on $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves and is founded on the hardness of the Multi-Target Group Action Inverse Problem (MT-GAIP). $\mathsf{CSI\text{-}SDVS}$ achieves strong security guarantees—Strong Unforgeability under Chosen-Message Attacks (SUF-CMA), Non-Transferability (NT), and Privacy of Signer’s Identity (PSI)—in the random oracle model, thereby making it among the most compact PQC-based SDVS schemes and the only post-quantum secure construction based on isogenies.
@@ 1.1 Related Work
Post-Quantum Constructions. In response to the vulnerabilities of classical assumptions, post-quantum SDVS schemes have been developed based on quantum-resistant foundations. Code-based SDVS constructions were proposed by Assar [8], Ren et al. [9], and Shooshtari et al. [10], leveraging the hardness of the bounded syndrome decoding problem in coding theory [11]. While these schemes are believed to be quantum-resistant, it was shown in [12] that the constructions by Ren et al. and Shooshtari et al. fail to satisfy the critical non-transferability property, which is central to the SDVS notion. One of the earliest lattice-based SDVS constructions was introduced by Wang et al. [14], employing preimage sampleable functions (PSFs) and Bonsai Trees for trapdoor delegation. Their work builds on the influential framework of Gentry et al. [13], which introduced PSFs as a core primitive in lattice-based signatures. Although the construction offers provable security in the random oracle model, it suffers from efficiency issues due to large key and signature sizes. Later, in 2016, Noh et al. [15] proposed a scheme in the standard model based on an LWE-based encryption scheme and a lattice-based chameleon hash function, albeit at the cost of significant computational overhead. To improve practicality, Cai et al. [16] introduced a more efficient scheme based on the Ring-SIS assumption, using rejection sampling to reduce signature size while maintaining security. More recently, Zhang et al. [17] proposed an SDVS construction combining SIS and LWE assumptions, further contributing to the development of quantum-secure SDVS protocols. However, lattice-based schemes generally incur key and signature sizes with complexity $\mathcal{O}(\lambda^{2})$ [14, 15, 16, 17], which may limit their suitability for resource-constrained environments.
@@ 1.2 Our Contribution
  • We present CSI-SDVS, an efficient post-quantum SDVS scheme that builds on ideal class group action on $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves. The security of our scheme is based on the hardness of the Multi-Target Group Action Inverse Problem (MT-GAIP) [24], and its design draws inspiration from the CSIDH [23] framework and the cyclic structure of the ideal class group as described in CSI-FiSh [24].
@@
  • Our isogeny-based construction naturally yields compact keys and signatures—a critical advantage over lattice-based SDVS schemes that inherently suffer from large key and signature sizes.
@@
Acknowledgements. I would like to express my gratitude to Sarah Arpin, Jason LeGrow, Frederik Vercauteren, and the Isocrypt Study Group for their constructive feedback, which helped improve this paper. I am also especially grateful to Péter Kutas for his insightful discussions during the early stages of the draft.
Outline. In the next section, we recall the formal definition of an SDVS scheme and provide the associated security requirements that such a scheme must satisfy. We then present a brief overview of the necessary mathematical prerequisites and describe the underlying group action-based hardness assumptions upon which our construction relies. In Section 3, we introduce our proposed SDVS scheme, CSI-SDVS. Section 4 analyzes the efficiency of our construction, and Section 5 provides a formal security proof.
@@ 2.1 Strong Designated Verifier Signature (SDVS) Scheme
[…]
@@ 2.2 Security Model
Query Phase: In this phase, $\mathcal{C}$ responds to polynomially many adaptive queries made by $\mathcal{A}$ by following the steps described below: Signing Oracle $\mathcal{O}_{\mathsf{Sign}}$: On receiving queries on a message $m$ from $\mathcal{A}$, $\mathcal{C}$ runs $\textsf{Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m)$ to obtain a signature $\sigma$ and sends it to $\mathcal{A}$. It then sets $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$. Simulating Oracle $\mathcal{O}_{\mathsf{Simul}}$: On receiving queries on a message $m$ from $\mathcal{A}$, $\mathcal{C}$ runs $\textsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m)$ to obtain a signature $\sigma$ and sends it to $\mathcal{A}$. It then sets $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$. Verifying Oracle $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving a message-signature pair $(m,\sigma)$ adaptively from $\mathcal{A}$, $\mathcal{C}$ runs $\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},\sigma,m)$ to obtain a decisional value $1$ for valid and 0 otherwise, and sends it to $\mathcal{A}$. Output: Eventually, $\mathcal{A}$ outputs a forgery $(m^{\ast},\sigma^{\ast})$, and wins if the following conditions hold: $\textsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},\sigma^{\ast},m^{\ast})=1.$ $(m^{\ast},\sigma^{\ast})\not\in\mathcal{Q}.$
@@
In an SDVS scheme, the notion of non-transferability (NT) ensures that the designated verifier cannot convince any third party of the authenticity of a valid signature. This is achieved by enabling the verifier to use the Simul algorithm to generate simulated signatures that are computationally indistinguishable from genuine ones produced by the signer. As a result, any conviction about the validity of a signature cannot be reliably transferred to others. The NT security of our SDVS construction is formalized as follows.
@@ 2.3 Elliptic Curves and Isogenies
Let $k:=\mathbb{F}_{q}$ denote a finite field, where $q:=p^{n}$ for a prime $p>3$ and a positive integer $n$. An elliptic curve $E$ over $k$ is a smooth, projective, genus-1 curve defined over $k$ with a distinguished point $\mathcal{O}_{E}$ serving as the identity element for the group law on $E$. For a positive integer $\ell$, the $\ell$-torsion subgroup of $E$ is defined as $E[\ell]:=\{P\in E(\overline{k})\mid[\ell]P=\mathcal{O}_{E}\}$. An elliptic curve is called supersingular if its group of $p$-torsion points is trivial over $\overline{\mathbb{F}}_{p}$, i.e., $E[p]=\{\mathcal{O}_{E}\}$. In particular, if $E/\mathbb{F}_{p}$ is supersingular, then $E(\mathbb{F}_{p})=p+1$.
@@
An endomorphism is an isogeny from an elliptic curve $E$ to itself. Examples include the multiplication-by-$m$ map $m]:P\mapsto[m]P$ for $m\in\mathbb{Z}$, and the Frobenius endomorphism $\pi:(x,y)\mapsto(x^{q},y^{q})$ for an elliptic curve defined over $\mathbb{F}_{q}$. The set of all endomorphisms of $E$ forms a ring under addition and composition, known as the endomorphism ring of $E$ and denoted by $\mathrm{End}(E)$. For a supersingular elliptic curve $E$ defined over $\mathbb{F}_{p^{2}}$, the $\mathrm{End}(E)$ is isomorphic to a maximal order in a quaternion algebra that is ramified at $p$ and $\infty$. In contrast, for a supersingular elliptic curve defined over $\mathbb{F}_{p}$, the ring of $\mathbb{F}_{p}$-rational endomorphisms, denoted by $\mathrm{End}_{\mathbb{F}_{p}}(E)$, is isomorphic to an order in some imaginary quadratic field. Hence, we have a strict inclusion $\mathrm{End}_{\mathbb{F}_{p}}(E)\subset\mathrm{End}(E)$.
@@
Notation. Following [24], if $\mathrm{Cl}(\mathfrak{O})$ is cyclic with generator $\mathfrak{g}$, then the map $\mathbb{Z}/N\mathbb{Z}\xhookrightarrow{}\mathrm{Cl}(\mathfrak{O})$ defined by $a\mapsto\mathfrak{g}^{a}$ gives a convenient representation of class group elements. Here, $N=\#\mathrm{Cl}(\mathfrak{O})$, and any element $\mathfrak{a}\in\mathrm{Cl}(\mathfrak{O})$ can be written as $\mathfrak{g}^{a}$ for some $a\in\mathbb{Z}/N\mathbb{Z}$. Using the shorthand $[a]$ for $\mathfrak{g}^{a}$ and $a]E:=\mathfrak{g}^{a}*E$, it follows that $[a][b]E=[a+b]E$.
@@ 2.4 Hardness Assumptions
(Group Operations) Given $g,h\in G$, compute $g^{-1}$, $gh$, and decide if $g=h$.
@@ 3 The Scheme: CSI-SDVS
In this section, we present CSI-SDVS, a new isogeny-based SDVS scheme. The construction is based on the CSIDH [23] framework and adopts the known ideal class group structure introduced in the CSI-FiSh signature scheme [24]. Specifically, CSI-FiSh exploits the precomputed structure of the ideal class group $\mathrm{Cl}(\mathfrak{O})$ for the CSIDH-512 parameter set, represented as a relation lattice of low-norm generators. In this setting, $\mathrm{Cl}(\mathfrak{O})$ is cyclic of order $N$, allowing each element to be uniquely expressed as $\mathfrak{g}^{a}]$, where $\mathfrak{g}$ is a fixed generator and $a\in\mathbb{Z}_{N}$. By leveraging this explicit cyclic structure, both deterministic computation and uniform sampling—achieved by exponentiating the fixed generator $\mathfrak{g}$ in $\mathbb{Z}_{N}$—are naturally supported, preserving the post-quantum security guarantees provided by the class group action framework. Consequently, CSI-SDVS achieves efficient and verifiable designated verifier signatures without relying on heuristic assumptions.
In CSI-SDVS, the key generation procedures for both the signer and the designated verifier closely follow the key generation phase of CSI-FiSh. The core of the $\mathsf{Sign}$ procedure—also mirrored in $\mathsf{Simul}$—is a structured interaction between class group elements and $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves. The signer begins with an intermediate curve $\hat{E}=[v]E_{0}$, samples a secret $b\in\mathbb{Z}_{N}$, and computes the target curve $Y=[b]\hat{E}$ via the class group action. The designated verifier, using its secret key $v\in\mathbb{Z}_{N}$, then computes $Y=[v]\bar{E}=[v+b]E_{0}$ from the intermediate curve $\bar{E}=[b]E_{0}$. This resulting curve constitutes the central component for verifying the designated verifier signature.
@@
A trusted party executes the setup procedure to generate the public parameters $\mathsf{pp}$ as follows. First, it selects a large prime $p=4\ell_{1}\ell_{2}\dots\ell_{n}-1$, where the $\ell_{i}$ are small, distinct, odd primes. Specifically, it sets $n=74$, $\ell_{1}=3$, $\ell_{73}=373$, and $\ell_{74}=587$. Next, it fixes a base curve $E_{0}:y^{2}=x^{3}+x$ defined over $\mathbb{F}_{p}$, such that $E_{0}\in\mathcal{E}ll_{p}(\mathfrak{O})$, and chooses a generator $\mathfrak{g}$ of the class group $\mathcal{G}=\text{Cl}(\mathfrak{O})$ with class number $N\approx p^{1/2}$, where $\mathfrak{O}=\mathbb{Z}[\sqrt{-p}]$. It then samples a cryptographically secure hash function $\mathcal{H}:\{0,1\}^{\ast}\rightarrow\{0,1\}^{\lambda}$, where $\lambda$ is a security parameter. In addition, a positive integer $\eta$ is specified. Finally, it outputs the public parameters $\mathsf{pp}:=(p,\mathfrak{g},N,E_{0},\mathcal{H},\eta)$.
@@
    Compute the curves $Y^{\prime}_{i}:=[y_{i}]E_{i}$ for each $i\in[\eta]$.